OpenStudy (anonymous):

f(x)=[(x+3)/2][x^2-4x+9] I need to find its derivative using product rule , Please Help !

OpenStudy (amistre64):

then use the product rule ... whats the hold up?

OpenStudy (anonymous):

how can i find its derivative ?

OpenStudy (amistre64):

what is the product rule?

OpenStudy (anonymous):

u'(x).v(x) + v'(x)u(x)

OpenStudy (anonymous):

u'(x).v(x) + v'(x).u(x)

OpenStudy (amistre64):

looks right so lets just use that you have pretty simple u and v

OpenStudy (amistre64):

whats your attempt look like?

OpenStudy (anonymous):

Attempt means? final answer?

OpenStudy (amistre64):

attempt means, you need help but you know the product rule so youve at least tried to work the problem. lets see what youve tried

OpenStudy (anonymous):

U'= 1/2 V'=2x-4 =(1/2)(x2-4x+9) + (2x-4)(x+3)/2

OpenStudy (anonymous):

I simplified it and got the simplification as 3x^2 /2 - 4x - 3/2

OpenStudy (anonymous):

I got 3x^2-3 all over 2. Check my work though for mistakes: (1/2)(x2-4x+9) + (2x-4)(x+3)/2 (1/2)(x2-4x+9) + (2x-4)(x/2+3/2) x^2/2-4x/2+9/2+2x^2/2-4x/2+6x/2-12/2 x^2/2-2x+9/2+x^2+2x-12/2 x^2/2+2x^2/2-3/2 2x^2+x^2-3 3x^2-3 all over 2 OR 3x^2/2-3/2 Hope this helps, let me know if you see any errors in my work.

OpenStudy (anonymous):

Thanks :)